OpenStudy (anonymous):

what is the secant of theta= 5pi/4?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

i know secant is 1/cos the cosine is -root2/2

OpenStudy (anonymous):

\[\sec \theta=5\pi/4\] If sec (Arccos) is the inverse of cos, then it follows that \[\theta = \cos 5\pi/4\]

OpenStudy (anonymous):

1/ neg root2/2

OpenStudy (anonymous):

right, so if : \[\cos(\theta) = -\frac{\sqrt{2}}{2}\] then we have: \[\frac{1}{\cos(\theta)} = \sec(\theta) = -\frac{1}{\frac{\sqrt{2}}{2}} = -\frac{2}{\sqrt{2}} = -\sqrt{2}\]

OpenStudy (anonymous):

I'm not sure what you're trying to say, but if I had to take a guess I'd say you're talking about the value of cos for an angle on a unit circle. Don't know why you're bringing it up though

OpenStudy (anonymous):

i was under the impression that you wanted to know: \[\sec(\frac{5\pi}{4})\]thats what i sounded like from the chat.

OpenStudy (anonymous):

No, she wants to find: \sec \theta = 5\pi/4\]

OpenStudy (anonymous):

\[\sec \theta = 5\pi/4\]

OpenStudy (anonymous):

Just use your calculator to find cos5π/4